# Quasi-Bayes empirical Bayes: a sequential approach to the Poisson compound decision problem

### v1

@@ F.1 Notation
$\liminf_{n}E^{*}\left(\sum_{i=1}^{D}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n,i}}\mathbf{h}_{i}(\mathbf{g}_{n})\right)=\liminf_{n}E^{*}\left(\sum_{i=1}^{D}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n,i}}\left(\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}\frac{k_{\mathbf{y}}({\bm{\vartheta}}_{i})\mathbf{g}_{n,i}p_{\mathbf{g}^{*}}(\mathbf{y})}{p_{\mathbf{g}_{n}}(\mathbf{y})}-\mathbf{g}_{n,i}\right)\right)$
$=\liminf_{n}E^{*}\left(\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}\frac{p_{\mathbf{g}^{*}}(\mathbf{y})^{2}}{p_{\mathbf{g}_{n}}(\mathbf{y})}-1\right)$
$>\liminf_{n}E^{*}\left(\left(\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}\frac{p_{\mathbf{g}_{n}}(\mathbf{y})}{p_{\mathbf{g}^{*}}(\mathbf{y})}p_{\mathbf{g}^{*}}(\mathbf{y})\right)^{-1}-1\right)=0,$
where the strict inequality holds since, under the contradiction assumption, $\mathds{P}^{*}(p_{\mathbf{g}_{n}}\not\rightarrow p_{\mathbf{g}^{*}})>0$. Thus there exists a constant $c>0$ such that
$E^{*}\left(-\sum_{i=1}^{D}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n,i}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})\right)<-c\alpha_{n+1}$
for $n$ large enough. On the other hand, denoting by $E^{*}_{n}$ the conditional expectation given $\mathcal{G}_{n}$, we can write that
$E^{*}\left(\sum_{i=1}^{D}\frac{\mathbf{g}^{*}_{i}}{(\tilde{\mathbf{g}}_{n,i})^{2}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})^{2}\right)$
$\leq E^{*}\left(\sum_{i=1}^{D}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n,i}^{2}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})^{2}\right)+E^{*}\left(\sum_{i=1}^{D}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n+1,i}^{2}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})^{2}\right)$
$\leq\alpha_{n+1}^{2}E^{*}\left(\sum_{i=1}^{D}\mathbf{g}^{*}_{i}\left(\frac{k_{\mathbf{Y}_{n+1}}({\bm{\vartheta}}_{i})}{p_{\mathbf{g}_{n}}(\mathbf{Y}_{n+1})}-1\right)^{2}\right)+\alpha_{n+1}^{2}E^{*}\left(\sum_{i=1}^{D}\mathbf{g}^{*}_{i}\frac{\left(\frac{k_{\mathbf{Y}_{n+1}}({\bm{\vartheta}}_{i})}{p_{\mathbf{g}_{n}}(\mathbf{Y}_{n+1})}-1\right)^{2}}{\left(1+\alpha_{n+1}\left(\frac{k_{\mathbf{Y}_{n+1}}({\bm{\vartheta}}_{i})}{p_{\mathbf{g}_{n}}(\mathbf{Y}_{n+1})}-1\right)\right)^{2}}\right)$
$\leq\alpha_{n+1}^{2}E^{*}\left(1+E^{*}_{n}\left(\sum_{i=1}^{D}\left(\frac{k_{\mathbf{Y}_{n+1}}({\bm{\vartheta}}_{i})}{\sum_{i=1}^{D}k_{\mathbf{Y}_{n+1}}({\bm{\vartheta}}_{i})\mathbf{g}_{n,i}}\right)^{2}\right)\right)$
$\quad\quad+\frac{\alpha_{n+1}^{2}}{(1-\alpha_{n+1})^{2}}E^{*}\left(1+E^{*}_{n}\left(\sum_{i=1}^{D}\left(\frac{k_{\mathbf{Y}_{n+1}}({\bm{\vartheta}}_{i})}{\sum_{i=1}^{D}k_{\mathbf{Y}_{n+1}}({\bm{\vartheta}}_{i})\mathbf{g}_{n,i}}\right)^{2}\right)\right)$
@@
$E^{*}\left(\mathcal{L}(\mathbf{g}_{n+1})\right)\leq E^{*}\left(\mathcal{L}(\mathbf{g}_{n})\right),$
for $n$ large enough, which contradicts $E^{*}(\mathcal{L}(\mathbf{g}_{n}))\rightarrow\infty$. ∎
@@
$||\mathbf{g}_{n}-\mathbf{g}^{*}||^{2}=o(n^{-\delta})\quad\mathds{P}^{*}-a.s.$
@@
In this proof, $\mathbf{g}_{n}$ and $\mathbf{g}^{*}$ are considered as random vectors in $\{\mathbf{v}\in\mathbb{R}^{D}:||\mathbf{v}||\leq 1\}$. By Lemma F.4, $\mathbf{g}_{n}$, seen as vectors in $\Delta^{D}$ converge $P^{*}$-a.s. to $\mathbf{g}\in\Delta^{D}$. This implies that $\mathbf{g}_{n}$ converges to $\mathbf{g}^{*}$ also when seen as vectors in $\mathbb{R}^{D}$. Moreover, $(\mathbf{g}_{n})$ satisfies the stochastic approximation in $\mathbb{R}^{D}$:
@@
  • B2
    $\sum_{n=1}^{\infty}\alpha_{n}^{1-\delta/2}{\bm{\epsilon}}_{n}$ converges, $\mathds{P}^{*}$-a.s.
  • B3
@@
Theorem 3.1.1 in Chen (2005) also requests the following condition: B4 There exists a is continuously differentiable function $\mathscr{l}:\{\mathbf{v}\in\mathbb{R}^{D}:||\mathbf{v}||\leq 1\}\rightarrow\mathbb{R}$ such that $\sup_{\delta_{1}\leq{\rm d}(\mathbf{g},\mathbf{g}^{*})\leq\delta_{2}}\mathbf{h}(\mathbf{g})^{T}\nabla_{\mathbf{g}}\mathscr{l}<0$ for every $\delta_{2}>\delta_{1}>0$. However, a close inspection of the proof of Theorem 3.1.1 in Chen (2005) shows that B4 is employed only to ensure convergence of $\mathbf{g}_{n}$ towards $\mathbf{g}^{*}$. In our case, this property has been proved in Lemma F.4.
B1 is obvious. To prove B2, notice that $(\sum_{k=1}^{n}\alpha_{k}^{1-\delta/2}{\bm{\epsilon}}_{k})_{n\geq 1}$ is a martingale. Moreover, there exoists a constant $C$ such that, for every $i=1,\dots,D$,
$\sup_{n}E^{*}((\sum_{k=1}^{n}\alpha_{k}^{1-\delta/2}{\bm{\epsilon}}_{k,i})^{2})$ $=\sup_{n}\sum_{k=1}^{n}\alpha_{k}^{2-\delta}E^{*}({\bm{\epsilon}}_{k,i}^{2})$
@@
since $\gamma(2-\delta)>1$ and $({\bm{\epsilon}}_{n,i})$ are uniformly bounded. We now prove B3. The function $\mathbf{h}$ is continuously differentiable at $\mathbf{g}^{*}$ and

### v2

@@ F.1 Notation
$\liminf_{n}E_{G^{*}}\left(\sum_{i=1}^{D}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n,i}}\mathbf{h}_{i}(\mathbf{g}_{n})\right)=\liminf_{n}E_{G^{*}}\left(\sum_{i=1}^{D}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n,i}}\left(\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}\frac{k_{\mathbf{y}}({\bm{\vartheta}}_{i})\mathbf{g}_{n,i}p_{\mathbf{g}^{*}}(\mathbf{y})}{p_{\mathbf{g}_{n}}(\mathbf{y})}-\mathbf{g}_{n,i}\right)\right)$
$=\liminf_{n}E_{G^{*}}\left(\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}\frac{p_{\mathbf{g}^{*}}(\mathbf{y})^{2}}{p_{\mathbf{g}_{n}}(\mathbf{y})}-1\right)$
$>\liminf_{n}E_{G^{*}}\left(\left(\sum_{\mathbf{y}\in\mathbb{N}_{0}^{k}}\frac{p_{\mathbf{g}_{n}}(\mathbf{y})}{p_{\mathbf{g}^{*}}(\mathbf{y})}p_{\mathbf{g}^{*}}(\mathbf{y})\right)^{-1}-1\right)=0,$
where the strict inequality holds since, under the contradiction assumption, $\mathds{P}_{G^{*}}(p_{\mathbf{g}_{n}}\not\rightarrow p_{\mathbf{g}^{*}})>0$. Thus there exists a constant $c>0$ such that
$E_{G^{*}}\left(-\sum_{i=1}^{D}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n,i}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})\right)<-c\alpha_{n+1}$
for $n$ large enough. On the other hand, denoting by $\mathcal{G}_{n}$ the sigma-algebra generated by $\mathbf{Y}_{1},\dots,\mathbf{Y}_{n}$, we can write that
$E_{G^{*}}\left(\sum_{i=1}^{D}\frac{\mathbf{g}^{*}_{i}}{(\tilde{\mathbf{g}}_{n,i})^{2}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})^{2}\right)$
$\leq E_{G^{*}}\left(\sum_{i=1}^{D}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n,i}^{2}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})^{2}\right)+E_{G^{*}}\left(\sum_{i=1}^{D}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n+1,i}^{2}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})^{2}\right)$
$\leq\alpha_{n+1}^{2}E_{G^{*}}\left(\sum_{i=1}^{D}\mathbf{g}^{*}_{i}\left(\frac{k_{\mathbf{Y}_{n+1}}({\bm{\vartheta}}_{i})}{p_{\mathbf{g}_{n}}(\mathbf{Y}_{n+1})}-1\right)^{2}\right)+\alpha_{n+1}^{2}E_{G^{*}}\left(\sum_{i=1}^{D}\mathbf{g}^{*}_{i}\frac{\left(\frac{k_{\mathbf{Y}_{n+1}}({\bm{\vartheta}}_{i})}{p_{\mathbf{g}_{n}}(\mathbf{Y}_{n+1})}-1\right)^{2}}{\left(1+\alpha_{n+1}\left(\frac{k_{\mathbf{Y}_{n+1}}({\bm{\vartheta}}_{i})}{p_{\mathbf{g}_{n}}(\mathbf{Y}_{n+1})}-1\right)\right)^{2}}\right)$
$\leq\alpha_{n+1}^{2}E_{G^{*}}\left(1+E_{G^{*}}\left(\sum_{i=1}^{D}\left(\frac{k_{\mathbf{Y}_{n+1}}({\bm{\vartheta}}_{i})}{\sum_{i=1}^{D}k_{\mathbf{Y}_{n+1}}({\bm{\vartheta}}_{i})\mathbf{g}_{n,i}}\right)^{2}\mid\mathcal{G}_{n}\right)\right)$
$\quad\quad+\frac{\alpha_{n+1}^{2}}{(1-\alpha_{n+1})^{2}}E_{G^{*}}\left(1+E_{G^{*}}\left(\sum_{i=1}^{D}\left(\frac{k_{\mathbf{Y}_{n+1}}({\bm{\vartheta}}_{i})}{\sum_{i=1}^{D}k_{\mathbf{Y}_{n+1}}({\bm{\vartheta}}_{i})\mathbf{g}_{n,i}}\right)^{2}\mid\mathcal{G}_{n}\right)\right)$
@@
$E_{G^{*}}\left(\mathcal{L}(\mathbf{g}_{n+1})\right)\leq E_{G^{*}}\left(\mathcal{L}(\mathbf{g}_{n})\right),$
for $n$ large enough, which contradicts $E_{G^{*}}(\mathcal{L}(\mathbf{g}_{n}))\rightarrow\infty$. ∎
@@
$||\mathbf{g}_{n}-\mathbf{g}^{*}||^{2}=o(n^{-\delta})\quad\mathds{P}_{G^{*}}-a.s.$
@@
In this proof, $\mathbf{g}_{n}$ and $\mathbf{g}^{*}$ are considered as random vectors in $\{\mathbf{v}\in\mathbb{R}^{D}:||\mathbf{v}||\leq 1\}$. By Lemma F.4, $\mathbf{g}_{n}$, seen as vectors in $\Delta^{D}$ converge $P_{G^{*}}$-a.s. to $\mathbf{g}\in\Delta^{D}$. This implies that $\mathbf{g}_{n}$ converges to $\mathbf{g}^{*}$ also when seen as vectors in $\mathbb{R}^{D}$. Moreover, $(\mathbf{g}_{n})$ satisfies the stochastic approximation in $\mathbb{R}^{D}$:
@@
  • T2
    $\sum_{n=1}^{\infty}\alpha_{n}^{1-\delta/2}{\bm{\epsilon}}_{n}$ converges, $\mathds{P}_{G^{*}}$-a.s.
  • T3
@@
Theorem 3.1.1 in Chen (2005) also requests the following condition:
T4 There exists a is continuously differentiable function $\mathscr{l}:\{\mathbf{v}\in\mathbb{R}^{D}:||\mathbf{v}||\leq 1\}\rightarrow\mathbb{R}$ such that $\sup_{\delta_{1}\leq{\rm d}(\mathbf{g},\mathbf{g}^{*})\leq\delta_{2}}\mathbf{h}(\mathbf{g})^{T}\nabla_{\mathbf{g}}\mathscr{l}<0$ for every $\delta_{2}>\delta_{1}>0$. However, a close inspection of the proof of Theorem 3.1.1 in Chen (2005) shows that T4 is employed only to ensure convergence of $\mathbf{g}_{n}$ towards $\mathbf{g}^{*}$. In our case, this property has been proved in Lemma F.4.
B1 is obvious. To prove T2, notice that $(\sum_{k=1}^{n}\alpha_{k}^{1-\delta/2}{\bm{\epsilon}}_{k})_{n\geq 1}$ is a martingale. Moreover, there exoists a constant $C$ such that, for every $i=1,\dots,D$,
$\sup_{n}E_{G^{*}}((\sum_{k=1}^{n}\alpha_{k}^{1-\delta/2}{\bm{\epsilon}}_{k,i})^{2})$ $=\sup_{n}\sum_{k=1}^{n}\alpha_{k}^{2-\delta}E_{G^{*}}({\bm{\epsilon}}_{k,i}^{2})$
@@
since $\gamma(2-\delta)>1$ and $({\bm{\epsilon}}_{n,i})$ are uniformly bounded. We now prove T3. The function $\mathbf{h}$ is continuously differentiable at $\mathbf{g}^{*}$ and